OpenStudy (anonymous):

Directions: Without graphing, tell whether each point is on the graph of the given line. Problem: -4x+2y=8; (1, 6)

OpenStudy (anonymous):

Substitute the x-value given into the equation. Then, rearrange and solve for y. If you get the corresponding y-value, the point is therefor on the line.

OpenStudy (anonymous):

so i +4x

OpenStudy (anonymous):

to both sides

OpenStudy (anonymous):

-4(1)+2y=8. Solve that for y.

OpenStudy (anonymous):

Add 4 to both sides and divide by 2.

OpenStudy (anonymous):

oh then i would get 2y=8+4x

OpenStudy (anonymous):

Yes. Now put in your x-value and see if the y-value given is produced.

OpenStudy (anonymous):

how do i do that?

OpenStudy (anonymous):

odnt i still have to get the y by itself

OpenStudy (anonymous):

dont*

OpenStudy (anonymous):

Well, you have 2y=8+4x. You're given the point (1,6). So, put in the x and y-values to see if they work: 2(6)=8+4(1). And it sure does!

OpenStudy (anonymous):

thank you! you helped me so much! xD

OpenStudy (anonymous):

No problem! :)